OpenStudy (anonymous):

Solve. x2 – 9x = –14 Answer A. x = –2 and x = 7 B. x = 2 and x = –7 C. x = –2 and x = –7 D. x = 2 and x = 7

OpenStudy (anonymous):

** x^2 - 9x = -14

OpenStudy (diyadiya):

D

OpenStudy (anonymous):

x^2 - 9x = -14 x^2 - 9x + 14 = 0 x^2 - 7x - 2x + 14 = 0 x(x - 7) - 2(x - 7) = 0 (x - 2)(x - 7) = 0 x - 2 = 0 ==> x = 2 x - 7 = 0 ==> x = 7 so, D. x = 2 and x = 7